OpenStudy (anonymous):

add and simplify. (3)/(y^2-3y+2)+(5)/(y^2-1)

jhonyy9 (jhonyy9):

3/(y-2)(y-1) +5/(y-1)(y+1) =((3(y+1)+5(y-2))/((y-2)(y-1)(y+1))

jhonyy9 (jhonyy9):

=(3y+3+5y-10)/((y-2)(y-1)(y+1))=(8y-7)/((y-2)(y-1)(y+1))

jhonyy9 (jhonyy9):

ok first you need factorizing y2 -3y +2 so for this you need getting two numbers so that their sum being equal -3 and their product being equal +2 so this two number is -2 and -1 so hence this will be (y-2)(y-1) so this is understandably ?

jhonyy9 (jhonyy9):

and for y2 -1 you can us formula (a2 -b2)=(a-b)(a+b) than will be y2 -1=(y-1)(y+1)

jhonyy9 (jhonyy9):

after this you need calcule the denominator common for those fractions ,yes ?

jhonyy9 (jhonyy9):

this denominator common you can getting if you multiply those parantheses all together

jhonyy9 (jhonyy9):

so after you need multiply those numerators with those parts of parantheses od denominator common with what you see that is necessary

OpenStudy (anonymous):

oh okay thanks:)